OpenStudy (anonymous):

7th grade math help please

OpenStudy (anonymous):

@awesomeness812

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

C. 4y+3y=28 y=4 This one's right. ------------------ D. 5w+6w-3w=64 w=8 Wrong ------------- E. t-2t-3t=32 t=8 Wrong

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

so it is a??????

hartnn (hartnn):